1.  An archeologist found the remains of an ancient wheel, which she then placed on a grid. If an arc of the wheel passes through A(-7,0), B(-3,4) and C(7,0), located the center of the wheel, and the standard equation of the circle defining its boundary. 


1
Expert's answer
2022-01-20T10:53:16-0500

The standard equation of the circle defining its boundary is



"(x-a)^2 + (y-b)^2 = R^2"

where the point with the coordinates (a, b) is the center of the circle. Put each point in the equation. Then we get the system of equations:



"\left\{
\begin{array}{rcl}
{(-7-a)^2 + (0 - b)^2 = R^2,} \\
{(-3-a)^2 + (4 - b)^2 = R^2,} \\
{(7-a)^2 + (0 - b)^2 = R^2.}
\end{array}
\right."

Open brackets (use the formula "(a+b)^2 = a^2 + b^2 + 2ab") :



"\left\{
\begin{array}{rcl}
{49 +a^2+ 14a + b^2 = R^2,} \\
{9+a^2 +6a+ 16+ b^2 - 8b = R^2,} \\
{49 +a^2- 14a + b^2 = R^2.}
\end{array}
\right."

Subtract from the first equation the third:

"49 +a^2+ 14a + b^2 - (49 +a^2- 14a + b^2) = R^2 -R^2"

"28a = 0 \implies a = 0"

Equate the first equation to the second:

"49 +a^2+ 14a + b^2 = 9+a^2 +6a+ 16+ b^2 - 8b"

Because "a = 0":

"49 +b^2 = 9+16+ b^2 - 8b"

"25 - 8b = 49
-8b = 24 \implies b = -3"

Then the center of the circle is (0, -3).

Find the radius of the circle: put point and the coordinates of center in the equation:

"(-7-0)^2 + (0 - (-3))^2 = R^2"

"49 +9 = R^2 \implies R =\sqrt{58}"

The standard equation of the circle:



"x^2 + (y+3)^2 = 58"